OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -1/4 and passes through the point (8, -1).

OpenStudy (zepp):

http://openstudy.com/study#/updates/500465dae4b0fb991133cf80 Here you go :) Situation 3.

OpenStudy (anonymous):

this still aint makin sense

OpenStudy (zepp):

We have a point at \(\large(8,-1)\) and the slope is \(\large -\frac{1}{4}\)

OpenStudy (anonymous):

si

OpenStudy (zepp):

We'll use the point-slope form and plug everything in. \[\large (y-y_1)=m(x-x_1)\\\large y-(-1)=-\frac{1}{4}(x-8)\]

OpenStudy (zepp):

And you simply distribute! \[\large y+1=-\frac{1}{4}x+2\]

OpenStudy (zepp):

Slope-intercept has the form \(\large y=mx+b\) \[\large y=-\frac{1}{4}x+2-1\\ \large y=-\frac{1}{4}x+1\]

OpenStudy (anonymous):

thanks